OpenStudy (anonymous):

-> ......Need Help......<- Question 1 (Worth 1 points) (02.01) A social networking site recorded that 80% of the people who became members of the site one month were between 30 and 40 years old. The equation below represents this information, where x represents the total number of people who became members of the site that month. x = 0.8x + 72 How many people who became members of the site that month were between 30 and 40 years of age? 112 288 360 432

OpenStudy (ivancsc1996):

x-0.8x=72=0.2x x=72/0.2 0.8*72/0.2=288